OpenStudy (anonymous):

How do you express this in simplest form: sqrt4.5

OpenStudy (btaylor):

\[\sqrt{4.5}=\sqrt{9/2} = \sqrt{9}/\sqrt{2}=3/\sqrt{2}\] If you need to rationalize the denominator, multiply both the numerator and denominator by sqrt(2) to get \[(3\sqrt{2})/2\]

OpenStudy (anonymous):

Well, the answer is suppose to be: \[1.5\sqrt{2}\]

OpenStudy (anonymous):

Oh same thing! >.< Thank youu =D